OpenStudy (anonymous):

PLEASE HELP!! WILL MEDAL!!!! Which of the following is the explicit formula for the sequence below? 3, 11, 19, 27, ...

OpenStudy (anonymous):

A. an = 3 + 8(n - 1) B. an = 8 + 3n C. an = 3 + 8n D. an = 8 + 3(n - 1)